OpenStudy (anonymous):

Which statement is true about the product square root of 2(5 + square root of 8)? It is rational and equal to 7. It is rational and equal to 9. It is irrational and equal to 4 + square root of 10. It is irrational and equal to 4 + 5square root of 2.

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

\[\sqrt{2}(5+\sqrt{8}) \] That's the equation

OpenStudy (campbell_st):

well just distribute \[\sqrt{2}(5 + \sqrt{8} = 5 \times \sqrt{2} + \sqrt{2} \times \sqrt{8}\] you need to know a rule for radicals \[\sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\] hope it helps

OpenStudy (anonymous):

how do you know whats a or b

OpenStudy (anonymous):

?

OpenStudy (anonymous):

which one is it

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

a little help?

OpenStudy (campbell_st):

if you apply the rule to \[\sqrt{2} \times \sqrt{8} = \sqrt{ 2 \times 8}\]

OpenStudy (campbell_st):

you need to evaluate this and \[\sqrt{2} \times 5 = 5\sqrt{2}\] now combine you're answers with a +